# iMAD: Intelligent Multi-Agent Debate for Efficient and Accurate LLM Inference

### v1

@@ A.2 Extracted Features for Debate Decision Classification
To provide more details on the feature set, we summarize all extracted 41 features and their descriptions in Table 5, organized into five categories: (i) surface-level statistics (3 features from the input question only) such as token counts, which quantify response verbosity and completeness, as shorter or underdeveloped responses often correlate with lower reasoning quality; (ii) readability metrics (the input question, initial reasoning, and self-critique each with 2 features) like Flesch Reading Ease (davenport2014readabilitytweetsgeographiccorrelation) and Coleman-Liau Index (coleman), which assess how easy or complex a response is to read and potentially reveal vague or oversimplified reasoning; (iii) syntactic features (the input question, initial reasoning, and self-critique each with 1 features) such as parse tree depth, which indicate grammatical complexity and elaboration in reasoning structure; (iv) part-of-speech counts (the input question, initial reasoning, and self-critique each with 3 features) including noun, verb, and adjective frequencies, which help characterize the factual density, action-oriented phrasing, and descriptive richness of the response; and (v) uncertainty-related lexical cues (11 features from the input question, 4 features from the initial reasoning, and 5 features from the self-critique) that reflect reasoning quality, such as hedging phrases (e.g., “maybe”), certainty expressions (e.g., “definitely”), and contrastive connectives (e.g., “however” and “but”), which are commonly associated with hesitation or internal conflict in reasoning.
These features capture fine-grained characteristics of the response that relate to model confidence, uncertainty, and reasoning behavior, enabling the classifier to determine when MAD is likely to be beneficial based on the single-agent output. Since all features are derived directly from this structured output without requiring additional queries or repeated sampling, the classifier can make efficient and accurate decisions while preserving scalability and improving token efficiency.
@@
To understand which signals drive the debate decision, we evaluate two complementary metrics on the 41 features. SHAP importance (SHAP) quantifies each feature’s contribution to the classifier’s predicted score on each example. We report the mean absolute value of SHAP (typical range 0.01 to 0.5) in Fig. 3(a), where larger bars indicate stronger influence on decisions. PCA contributions (pca) capture global structure by aggregating feature loadings on principal components that explain most of the variance; larger bars indicate features that account for more variance (typical range 0.02 to 0.15 for each feature), as shown in Fig. 3(b). In both figures, signals from the self critique and the response structure rank near the top: uncertainty and conflict markers, length and brevity cues, and syntactic depth are consistently among the most influential features. The consistency between SHAP and PCA suggests that the classifier relies on stable, human understandable cues.
@@
To further investigate whether low-ranked features are still useful, we normalize the absolute SHAP values and PCA contributions for each feature, sum them to obtain a joint importance score for every feature, exclude the 20% of features with the lowest joint scores, and rerun the classifier. Concretely, we drop these eight features that consistently appear in the bottom 20%: qtype_do, qtype_who, qtype_when, qtype_how, qtype_where, qtype_is, qtype_does, and NegativeReason_CertaintyCount. Table 6 reports accuracy and tokens per dataset for iMAD with all features compared to the set that excludes these least important 20% features. As shown in Table 6, removing the low-ranked features leads to a small accuracy drop of about 0.5% and a noticeable increase in tokens of about 6.9% on average across six datasets, from 2,053 to 2,194 tokens per question. The higher token usage comes from less effective decisions that trigger debate on cases that are already correct, while missing a few recoverable hard cases mainly affects accuracy. These results justify that even features with lower importance scores in terms of SHAP and PCA rankings still contribute to more accurate debate decisions and help maintain token efficiency. Therefore, we use all 41 features.
@@
(b) PCA Contributions
Figure 3: Feature importance for the debate-decision classifier. Bars indicate mean absolute SHAP values (top) and aggregated PCA contributions across principal components that explain most variance (bottom).
Appendix B Additional Analysis
In this section, we provide additional performance analysis. In section B.1, we analyze accuracy per 100k total tokens (ApT). In section B.2, we analyze per-question latency and explain how the mix of input and output tokens affects total inference time.
B.1 Analysis on Accuracy per Token
To further evaluate token efficiency, we report accuracy per tokens (ApT), where tokens include both inputs and outputs per instance. For each method and dataset, we first measure the average number of token tokens per question, summing both input and output tokens. Given accuracy $Acc\in[0,1]$ and the average total token count $\text{Token}$ (see Table 3 for exact values), we define
$\mathrm{ApT}=Acc\cdot\frac{100,000}{\#~\text{Token}}.$
A larger ApT indicates better token efficiency, corresponding to more correct answers per 100,000 tokens of budget.
Table 7 summarizes ApT across all datasets. CoT attains the highest ApT because it uses very few tokens per question, but the Table 3 shows that its prediction accuracy is consistently lower than that of multi-agent methods. SC improves accuracy over CoT by running CoT multiple times and aggregating the outputs, but the additional generations make its ApT much smaller, since the token cost increases faster than the accuracy gain. Full-debate MAD baselines such as MAD and GD achieve higher accuracy but pay a high token cost, which leads to very low ApT. Selective debate methods strike a better balance. Both DOWN and iMAD obtain much higher ApT than MAD and GD because they skip debating on many questions that the single agent already answers correctly. Compared with DOWN, iMAD has slightly lower ApT, about $0.92\times$ on average across the six datasets, because it still triggers debates on questions where debate is necessary to improve the answer. These debates consume many tokens, so ApT becomes slightly lower even though accuracy is higher. This modest extra token cost yields an average accuracy gain of about 4.1% without any threshold tuning on the evaluation datasets. In practice, this tradeoff is often favorable: iMAD retains most of the token savings of selective debate while achieving higher accuracy consistently across datasets. Thus, Table 7 should be read together with the accuracy table Table 3. ApT shows that iMAD preserves strong token efficiency compared with full-debate MAD and single-agent baselines, while the accuracy results highlight its consistent gains over DOWN.
Dataset Single-Agent
Full-Debate
MAD
Selective
MAD
@@ B.1 Analysis on Accuracy per Token
MEDQA 117.3 22.2 20.3 4.8 68.2 63.1
MMLU 113.6 23.4 26.7 6.2 98.0 88.3
GSM8K 115.4 20.6 22.2 4.8 89.4 82.7
OKVQA 45.4 8.1 11.5 2.6 37.6 34.7
VQA-v2 34.5 5.5 9.2 2.3 24.1 23.3
ScienceQA 50.0 8.8 13.2 3.2 34.5 31.4
Average 79.4 14.8 17.2 4.0 58.6 53.9
Table 7: Accuracy per 100k tokens (ApT). ApT measures the expected number of correct answers obtained for every 100,000 tokens, so larger values indicate better token efficiency.
@@
Table 8: Inference time (s) per question comparison of iMAD and baselines across datasets.
@@ B.2 Inference Time Analysis
To evaluate the computational efficiency of iMAD, we measure the per-question inference time across different methods. As shown in Table 8, CoT achieves the lowest inference latency but yields lower accuracy. SC improves accuracy by running CoT multiple times and aggregating the outputs, but this significantly increases latency (up to 5.3–15.2s). Full-debate MAD frameworks such as MAD and GD incur even higher latency, with inference times reaching up to 49.2s per question. In contrast, selective MAD methods such as DOWN and iMAD maintain near single-agent latency, with inference times ranging from 1.1 to 1.8s. However, unlike DOWN, which suffers from lower accuracy due to its inability to generalize across different models or datasets, iMAD consistently matches or exceeds the accuracy of full-debate MAD frameworks. On average, iMAD reduces inference time by 32.6% and 68.9% compared to MAD and GD, respectively, while preserving low latency. It is worth noting that latency is not directly proportional to the total token cost. Table 3 reports total tokens (input and output), and Table 9 decomposes them into input and output (out). Output tokens generally take longer per token to generate than input tokens, although very large inputs still add measurable time. For example, GSM8K uses fewer total tokens than MMLU (618 vs 764), yet produces more output tokens on average (203 vs 154), so decoding takes longer (1.6s vs 1.0s). Within a dataset, SC on VQA-v2 uses fewer output tokens than MAD (802 vs 2,450) but a much larger input (13,211 vs 6,346) because it aggregates multiple single-agent runs, which explains its higher latency despite lower outputs. Overall, these findings highlight the ability of iMAD to balance accuracy and computational cost effectively, making it a practical choice for real-world deployments where both accuracy and inference speed are critical.
Appendix C Additional Results
@@
In this section, we present two ablation studies and a cross-LLM evaluation to assess the effectiveness of iMAD ’s two key design components and to gauge its generality. The ablations examine (i) the structured self-critique prompting, designed to enable interpretable and informative feature extraction for downstream debate decision classification (in section C.1); and (ii) the FocusCal loss, designed to guide the classifier in skipping debates that are unnecessary (i.e., when the single-agent answer is already correct or not recoverable by MAD) and triggering them when beneficial (i.e., when MAD is likely to correct an initially wrong answer) (in section C.2). In section C.3, We present cross-LLM results on GPT-5 nano and Qwen 3.0 models.
C.1 Ablation Studies: Effectiveness of Structured Self-Critique Prompting
@@
To evaluate the impact of prompt design, we compare structured self-critique prompt in iMAD, which asks the model to first provide reason step by step and then explicitly critique its own answer, against a standard CoT prompt without self-critique. For this study, we report (i) input-output token usage per data instance and (ii) accuracy of the answer (i.e., final answer correctness), to assess the overall impact of the prompt on the performance of system. As shown in Table 10, ours incorporating self-critique consistently improves accuracy across all six datasets. This improvement is particularly notable on complex reasoning tasks; for example, GSM8K shows a 7.2% accuracy gain. Although the self-critique prompt slightly increases token usage (e.g., from 1,222 to 1,300 on MEDQA), this cost is justified by the substantial accuracy improvement.
This accuracy improvement comes from the structured self-critique prompt’s ability to elicit more detailed and diverse reasoning, including counterarguments and comparative confidence judgments. These elements help reveal subtle indicators of uncertainty or conflicting logic in the model’s response, which are often missing from standard CoT outputs. As a result, the debate decision classifier can more effectively determine whether a debate is warranted. For instance, the structured prompt may generate divergent reasoning paths or express comparable confidence across competing answers. The classifier learns to interpret such cues as signs of uncertainty, indicating that debate may improve the final answer. In contrast, standard CoT prompts typically produce less informative responses, which makes it more difficult to identify cases where debate could be beneficial. By leveraging lexical and semantic features such as contrastive phrasing, shifts in certainty, and justification depth, the 41 extracted features help the classifier make more accurate debate decisions.

### v2

@@ A.2 Features for Debate Decision Classification
To provide more details on the feature set, we present all 41 extracted features and their descriptions in Table 5, organized into five categories: (i) surface-level statistics (3 features from the input question only), such as token counts, which quantify question length and specificity, since very short or underdeveloped questions often correlate with lower reasoning quality; (ii) readability metrics (6 features, consisting of 2 features for each of the input question, initial reasoning, and self-critique), including Flesch Reading Ease (davenport2014readabilitytweetsgeographiccorrelation) and the Coleman-Liau Index (coleman), which assess how easy or complex a response is to read and potentially reveal vague or oversimplified reasoning; (iii) syntactic features (3 features, consisting of one feature for each of the input question, initial reasoning, and self-critique), such as parse tree depth, which indicate grammatical complexity and the degree of elaboration in the reasoning structure; (iv) part-of-speech counts (9 features, consisting of 3 features for each of the input question, initial reasoning, and self-critique), including noun, verb, and adjective frequencies, which help characterize the factual density, action-oriented phrasing, and descriptive richness of the response; and (v) uncertainty-related lexical cues (20 features, consisting of 11 from the input question, 4 from the initial reasoning, and 5 from the self-critique), which reflect reasoning quality, such as hedging phrases (e.g., “maybe”), certainty expressions (e.g., “definitely”), and contrastive connectives (e.g., “however” and “but”), and are commonly associated with hesitation or internal conflict in reasoning.
These features together capture fine-grained characteristics of the response that relate to model confidence, uncertainty, and reasoning behavior, enabling the classifier to determine when MAD is likely to be beneficial based on the single-agent output. Since all features are derived directly from this structured output without requiring additional queries or repeated sampling, the classifier can make efficient and accurate decisions while preserving scalability and improving token efficiency.
@@
To understand which features drive the debate decision, we evaluate feature importance using two complementary metrics over all 41 features. First, we use SHapley Additive exPlanations (SHAP) importance (SHAP), a feature-attribution method that has been widely used in the NLP and LLM settings to interpret model behavior (horovicz-goldshmidt-2024-tokenshap; SHAP-LLM-2024). For each feature, SHAP importance measures how much that feature contributes to the classifier’s predicted debate-triggering score on each data instance. For each feature, we compute the mean absolute SHAP value (typically ranging from $1.9\times 10^{-5}$ to $5.8\times 10^{-5}$ in our setting) and report these values in Fig. 3(a), where longer bars indicate stronger influence on the classifier’s decisions. Second, contributions from Principal Component Analysis (PCA) (pca) summarize how much each feature influences the top principal components, which are the directions in which the feature values vary the most across examples. Concretely, we run PCA on the standardized feature matrix, sum the absolute loadings of each feature over the principal components that explain most of the variance, and plot these aggregated contributions (typically ranging from 0.02 to 0.15 per feature) in Fig. 3(b). Similar to SHAP, longer bars in PCA features account for higher variance in the feature space.
Across both metrics, features derived from the self-critique consistently rank among the most influential. In particular, features capturing uncertainty and conflict signals (e.g., NegativeReason HedgeCount, NegativeReason contrast, NegativeReason SyntacticDepth, and readability features) appear near the top in both SHAP importance and PCA contributions. In addition, several question-level features such as QuestionToken and the question-type indicators (e.g., qtype why and qtype how) also contribute noticeably. Intuitively, longer questions and open-ended types like “why” or “how” often correspond to harder problems, so combining these signals with the uncertainty patterns in the self-critique helps the classifier better identify cases where debate is likely to be beneficial. Overall, the SHAP and PCA results indicate that the classifier mainly relies on self-critique features. It also uses readability, uncertainty-related lexical cues, and syntactic features from the question and initial reasoning to decide when triggering debate is likely to be beneficial.
@@
To further investigate whether low-ranked features are still useful, we examine the impact of these features on the classifier. For each feature, we first compute a joint importance score by normalizing its absolute SHAP value and its PCA contribution and then summing these two normalized quantities. The SHAP term reflects how strongly the feature affects the classifier’s prediction on individual examples, while the PCA term reflects how much the feature contributes to variation across the entire dataset. We then sort all 41 features by this joint score and remove the bottom 20% of features. Specifically, we remove the following eight features that fall into the bottom 20% by joint score: qtype_do, qtype_who, qtype_when, qtype_how, qtype_where, qtype_is, qtype_does, and NegativeReason_CertaintyCount. We retrain the debate decision classifier using the remaining 33 features and compare it with the full model that uses all 41 features. For both versions, we report accuracy and the average token costs per question on each dataset.
As shown in Table 6, removing these low-ranked features results in a small but consistent accuracy drop of about 0.5% on average and a noticeable increase in token usage of about 6.9% across the six datasets (from 2,053 to 2,194 tokens per question). The higher token cost arises since the pruned classifier makes less precise decisions about when debate is needed, triggering debate more often on questions that are already answered correctly. The accuracy drop occurs since the pruned classifier also misses some hard questions where debate could have corrected the initial answer. These results indicate that even features with low SHAP and PCA scores still help the classifier make more accurate and token-efficient decisions. Therefore, we keep all 41 features in our final design.
@@
(b) Aggregated PCA Contributions
Figure 3: The feature importance of 41 features used by the debate decision classifier: (a) mean absolute SHAP values and (b) aggregated PCA contribution for each feature, showing how much that feature contributes to the principal components that explain most of the variance across the dataset.
Appendix B Additional Analyses
In this section, we provide additional performance analyses. In Section B.1, we analyze accuracy per 100k total tokens (ApT) to show the token efficiency. In Section B.2, we analyze per-question latency and explain how the mix of input and output tokens affects total inference time.
Dataset Single-Agent Full-Debate MAD Selective MAD
@@ Appendix B Additional Analyses
ApT Acc. (%) ApT Acc. (%) ApT Acc. (%) ApT Acc. (%) ApT Acc. (%) ApT Acc. (%)
MEDQA 117.3 76.6 22.2 77.3 20.3 81.9 4.8 80.2 68.2 79.2 63.1 82.0
MMLU 113.6 86.8 23.4 88.2 26.7 89.5 6.2 82.6 98.0 88.3 88.3 89.2
GSM8K 115.4 71.3 20.6 74.5 22.2 76.4 4.8 73.4 89.4 72.6 82.7 84.8
OKVQA 45.4 88.3 8.1 89.2 11.5 89.8 2.6 87.3 37.6 88.1 34.7 90.3
VQA-v2 34.5 77.5 5.5 77.6 9.2 81.0 2.3 81.3 24.1 78.6 23.3 81.3
ScienceQA 50.0 86.0 8.8 86.2 13.2 89.4 3.2 87.4 34.5 87.0 31.4 90.8
Average 79.4 81.1 14.8 82.2 17.2 84.7 4.0 82.0 58.6 82.3 53.9 86.4
Table 7: Accuracy per 100k tokens (ApT) and corresponding accuracy (Acc.) from Table 3. ApT measures the expected number of correct answers obtained for every 100k tokens, so larger values indicate better token efficiency.
B.1 Analysis on Accuracy per Token (ApT)
To further evaluate token efficiency, we report Accuracy per Tokens (ApT) as an auxiliary metric, following recent works that measure token efficiency in terms of ApT for reasoning LLMs (yan-etal-2025-long; bi2025cotxadaptiveframeworkcrossmodel). For each method and dataset, we measure the average total number of tokens (input + output) used per question. Given accuracy $\text{Acc}\in[0,1]$ and the average total number of tokens (i.e., $\text{Token}$ in Table 3), ApT is defined as:
$\mathrm{ApT}=\text{Acc}\times\frac{100,000}{\#~\text{Token}}.$
A larger ApT means that, on average, a method produces more correct answers per 100,000 tokens, but ApT does not capture the final accuracy level. Some methods that use very few tokens (e.g., CoT) can appear highly efficient under ApT while still performing much worse in accuracy than multi-agent methods. Therefore, ApT should be viewed only as a coarse indicator of token usage and must be interpreted together with accuracy. In practice, a method with slightly lower ApT but substantially higher accuracy is usually more desirable than a method that maximizes ApT alone.
In Table 7, we report ApT and the corresponding accuracy across all datasets. CoT attains the highest ApT as it uses very few tokens per question, but its accuracy is still lower than that of multi-agent methods. SC improves accuracy over CoT by running multiple CoT generations and aggregating the outputs, but these extra generations substantially increase the token cost. As a result, SC’s ApT becomes much smaller, since token usage grows faster than the accuracy gain. Full-debate MAD baselines such as MAD and GD achieve higher accuracy than single-agent methods but need a larger token cost, since they trigger debate for every question and repeatedly broadcast the growing transcript to all agents. This high cost leads to low ApT values.
Selective debate methods (e.g., DOWN and iMAD) strike a better balance between accuracy and token usage. Both DOWN and iMAD obtain much higher ApT than MAD and GD since they skip debating on many questions that the single agent already answers correctly. Compared with DOWN, iMAD has slightly lower ApT on average (53.9 vs. 58.6) across the six datasets. This is because iMAD still triggers debates on questions where debate is necessary to improve the answer, while DOWN skips some of these cases. These additional debates consume extra tokens, so ApT becomes slightly lower even though iMAD achieves higher accuracy. iMAD can be applied in a zero-shot setting, where it uses a fixed debate decision classifier without any dataset-specific tuning, while DOWN requires access to labeled evaluation data to tune its confidence threshold, violating the standard zero-shot assumption typically upheld by single-agent and MAD baselines. Under this stricter zero-shot protocol for iMAD, the modest extra token cost still yields an average accuracy gain of about 4.1% over DOWN.
In practice, this trade-off is often favorable: iMAD retains most of the token savings of selective debate while achieving higher accuracy consistently across datasets. ApT shows that iMAD preserves strong token efficiency compared with full-debate MAD and single-agent baselines, while the accuracy results highlight its consistent gains over DOWN.
@@ B.1 Analysis on Accuracy per Token (ApT)
Table 8: Inference time per question comparison of iMAD and baselines across datasets.
@@ B.2 Inference Time Analysis
To evaluate the computational efficiency of iMAD, we measure the per-question inference time for each method on each dataset. Here, inference time refers to the time spent to process a single question and produce an answer, including all calls to the underlying LLM. As shown in Table 8, CoT achieves the lowest inference time but yields lower accuracy compared with multi-agent methods. SC improves accuracy over CoT by running CoT multiple times and aggregating the outputs. However, these extra generations substantially increase the overall computation, leading to longer inference times (up to 5.3–15.2 s per question). Full-debate MAD frameworks such as MAD and GD have much longer inference time, reaching up to 49.2 s per question, since they trigger debate for every question and repeatedly resend a growing shared transcript to all agents. In contrast, selective MAD methods such as DOWN and iMAD keep the inference time close to that of a single agent, ranging from 1.1 s to 1.8 s per question. Unlike DOWN, which often achieves lower accuracy due to its limited ability to generalize across different models or datasets, iMAD consistently matches or exceeds the accuracy of full-debate MAD frameworks while keeping inference time close to single-agent baselines.
On average, iMAD reduces inference time by 32.6% compared with MAD and by 68.9% compared with GD, while still maintaining low per-question inference time. It is important to note that inference time is not directly proportional to the total number of tokens. Table 3 reports the total number of tokens (input and output), and Table 9 decomposes them into input and output tokens. While additional input tokens incur a slightly longer processing time, generating output tokens is typically more time-consuming since the model needs to generate them sequentially. For example, GSM8K uses fewer total tokens than MMLU (618 vs. 764 tokens), yet produces more output tokens on average (203 vs. 154 tokens). As a result, GSM8K has a longer inference time (1.6 s vs. 1.0 s). However, a very large number of input tokens can also dominate the inference time. For example, within the VQA-v2 dataset, SC uses fewer output tokens than MAD (802 vs. 2,450 tokens) but a much larger number of input tokens (13,211 vs. 6,346 tokens) as it aggregates multiple single-agent runs. This much larger number of input tokens is the reason why SC has a longer inference time than MAD despite generating fewer output tokens.
Overall, these findings highlight that iMAD provides a favorable balance between accuracy and computational cost. It achieves accuracy comparable to or better than full-debate MAD methods while keeping inference time close to that of single-agent baselines, making it a practical choice for real-world deployments where both accuracy and inference speed are critical.
Appendix C Additional Evaluation Results
@@
To further assess both the effectiveness of the two key design components in iMAD and its generality across different LLMs, we present two ablation studies and a cross-LLM evaluation. The ablation studies focus on (i) the structured self-critique prompting, which is designed to enable interpretable and informative feature extraction for downstream debate decision classification (Section C.1); and (ii) the FocusCal loss, which is designed to guide the classifier skip debates that are unnecessary (when the single-agent answer is already correct or not recoverable by triggering MAD) and trigger debates when they are likely to correct an initially wrong answer (Section C.2). Finally, we present the results of the cross-LLM evaluation in Section C.3.
C.1 Ablation Study on Structured Self-Critique Prompt Design
@@
To evaluate the impact of prompt design, we compare the structured self-critique prompt used in iMAD with a standard CoT prompt that does not include self-critique. In the structured prompt, the model is asked to (i) provide an initial answer with step-by-step reasoning; and (ii) explicitly critique its own answer before committing to a final choice. In the standard CoT prompt, the model only produces step-by-step reasoning for a single answer without any self-critique. For this study, we report two quantities for each method and each dataset: (i) total (input + output) token usage per data instance, and (ii) accuracy of the answer (i.e., final answer correctness). These two metrics allow us to see how the self-critique prompt affects both token usage and accuracy performance. As shown in Table 10, the structured self-critique prompt in iMAD consistently improves accuracy across all datasets and tasks compared with the standard CoT prompt. The gain is more evident for complex reasoning tasks. For example, on math reasoning benchmarks such as GSM8K, the self-critique prompt improves accuracy by 7.2% over the standard CoT prompt. Although the structured self-critique prompt slightly increases token usage (e.g., from 1,222 to 1,300 tokens per question on MEDQA), this extra cost is small compared with the accuracy gain and is thus desirable in practice.
This accuracy improvement comes from the ability of the structured self-critique prompt to elicit more detailed and diverse reasoning. The prompt asks the model not only to justify its initial answer but also to generate counterarguments that challenge it. For example, an initial reasoning may say “The answer is option A because …,” while the self-critique then says “However, option B could also be correct because …” or “I am less certain about A after considering B.” These additional steps often reveal signals that are missing from standard CoT outputs, such as explicit mentions of alternative options, phrases that indicate doubt (e.g., “maybe”, “not sure”, and “could be”), or changes of mind between the initial answer and the final answer.
These signals help the debate decision classifier decide whether a debate is needed. For instance, if the self-critique proposes a new option and relies heavily on hedging phrases while comparing it with the original answer, the classifier can interpret this pattern as evidence of uncertainty, thereby triggering MAD. In contrast, if the self-critique repeats the same answer, gives a consistent justification, and contains few hedging words, the classifier is more likely to skip the debate. Standard CoT prompts usually produce only a single chain of reasoning that supports one answer, which makes it hard to distinguish questions that really need debate from those that do not.
@@
C.2 Ablation Study on FocusCal Loss
To evaluate the effectiveness of the proposed FocusCal loss in iMAD, we conduct ablation studies from three perspectives: (i) how each loss component (i.e., $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE) in FocusCal contributes to the final performance in terms of accuracy and token costs; (ii) how FocusCal outperforms standard loss designs, such as Binary Cross-Entropy (BCE) (li2024rediscoveringbcelossuniform) and Mean Squared Error (MSE) (koksoy2006multiresponse), on debate decision correctness; and (iii) what is the impact of the calibration term ECE in FocusCal by replacing ECE with BCE or MSE while keeping other terms of $L_{\text{AF}}$ and $L_{\text{CP}}$ fixed.
Loss Component Acc (%) # Token
$L_{\text{AF}}$ only 78.8 3,558
$L_{\text{CP}}$ only 78.1 3,379
ECE only 79.1 3,757
$L_{\text{AF}}$ + $L_{\text{CP}}$ 79.7 3,552
$L_{\text{AF}}$ + ECE 79.7 3,561
$L_{\text{CP}}$ + ECE 79.8 3,577
$L_{\text{AF}}$ + $L_{\text{CP}}$ + ECE (FocusCal) 81.3 3,489